An express and local train leave Gray’s Lake at 3 P.M. and head for Chicago 50 miles away. The express travels twice as fast as the local, and arrives 1 hour ahead of it. Find the speed of each train.
: An express and local train leave Gray’s Lake at 3 P.M. and head for Chicago 50 miles away. The express travels twice as fast as the local, and arrives 1 hour ahead of it. Find the speed of each train.

Answer by stanbon(18743)

An express and local train leave Gray’s Lake at 3 P.M. and head for Chicago 50 miles away. The express travels twice as fast as the local, and arrives 1 hour ahead of it. Find the speed of each train.
--------------------------------
Local train DATA:
rate = x mph ; distance = 50 miles ; time = 50/x hrs
----------------
Express train DATA:
rate = 2x mph ; distance = 50 miles ; time = 50/2x hrs
----------------
EQUATION:
local train time - express train time = 1 hr
50/x - 50/2x = 1
Multiply thru by 2x to get:
100 - 50 = 2x
2x = 50
x = 25 mph (local train speed)
2x = 50 mph (express train speed)
===================
Cheers,
Stan H.